40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 14 03:42:31 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
    Reply  Thu Jan 15 04:41:01 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
       Reply  Thu Jan 15 14:28:14 2015, Jenne, Update, LSC, LSC model change idea SwitchableErrorSignalProposal_Jan2015.pdf
          Reply  Fri Jan 16 03:31:35 2015, Jenne, Update, LSC, LSC model change implemented LSCorig_15Jan2015.pngLSCpost_15Jan2015.png
             Reply  Fri Jan 16 18:46:15 2015, Koji, Update, LSC, LSC model change implemented 
                Reply  Fri Jan 16 20:01:32 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Tue Jan 20 15:09:01 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Sat Mar 7 18:53:48 2015, Jenne, Update, LSC, DoF selector matrices replaced with filter banks DoFselectors_7March2015.png
Message ID: 10903     Entry time: Thu Jan 15 04:41:01 2015     In reply to: 10900     Reply to this: 10904
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Thoughts on going forward with variable finesse 

[Jenne, Diego]

Life would be easier with the UGF servos working.  As Diego already elogged, we aren't sure why the demod phases are changing, but that is certainly causing the I-signals to dip below zero, which the log function can't handle (there is a limiter before the log, so that the signal can't go below 1e-3).  Anyhow, this is causing the UGF servos to freak out, so we have not been using them for tonight's locking.

Our goal tonight was to see if we could introduce a nice big MICH offset, and then lower the CARM offset while keeping the arms locked on ALS.  We hope to see some kind of sign of a PDH signal in some RF PD. 

In the end, the highest we got to was -460 MICH offset counts, which we think is about 29nm (if our rough calibration is accurate). The MICH half fringe should be 188nm. With this offset, we got down to 0.3 CARM offset counts while locked on ALS.  We think that this is around 300pm, plus or minus a lot.  Note that while yesterday I had a pretty easy time getting to -660 counts of MICH offset, tonight I struggled to get past -200.  The only way we ended up getting farther was by lowering the CARM offset.  Although, as I type this, I realized that last night's work already had a lower CARM offset, so maybe that's key to being able to increase the MICH offset. 

We watched REFL11I and REFL11I/(TRX+TRY) on striptool, but we didn't see any evidence of a PDH signal.  We lost lock when I tried to transition CARM over to REFLDC, but I wasn't careful about my offset-setting, so I am not convinced that REFLDC is hopeless.

So.  Tonight, we didn't make any major locking progress (the MC started being fussy for about an hour, right after I ran the LSC offsets script, just before we started locking in earnest).  However, we have some ideas from talking with Rana about directions to go:

* Can we transition CARM over to REFL11I, and then engage the AO path?

* Then, while the MICH offset is still large, can we transition DARM over to POX or POY, actuating on a single arm?  If CARM is totally suppressed, this is DARM-y.  If CARM doesn't have the AO path yet, this is halfsy-halfsy, but maybe we don't care.

* Then, can we lower the MICH offset and transition back to a REFLQ signal?

* Separately, it seemed like we kept losing PRC lock due to PRC motion.  If the MICH offset is very large, are we sideband-limited at the POP port, such that we can use the POP DC QPD?  Is it even worth it?

 

MICH calibration:

A single mirror (ITM) moving by lambda/2, in the MICH-only situation is the full range, from bright to dark fringe.  So, half fringe should be lambda/4, or about 133nm.  If we are thinking about pushing on the BS, there's an extra factor of sqrt(2), so I think the half fringe should be at 188nm of BS motion.

When we had MICH locked on ASDC/POPDC, we put in a line at 143.125Hz, at 20 counts to (0.5*BS-0.2625*PRM), so a total of 10 counts to the BS at 143Hz.  Given the BS calibration in http://nodus.ligo.caltech.edu:8080/40m/8242, this is 10.1pm of actuation.  We saw a line in the error signal of 0.1 counts, so we infer that the MICH error signal of ASDC/POPDC has a calibration of 94pm/count. This number was invariant over a few different MICH offsets, although the ones I measured at were all below about 100 counts of MICH offset, so maybe this number changes as we start to get farther from the MICH dark fringe.

 

IFO left flashing (all mirrors aligned except SRM) in case anyone wants fresh data for that.

ELOG V3.1.3-